Determine the type of number that the solutions of:
x^2+8x-9=0 will be

Answer by stanbon(75887)

Determine the type of number that the solutions of:
x^2+8x-9=0 will be
---
discriminant = b^2-4ac = 8^2-4*1*-9 = 64+36 = 100
---
Since the discrim is positive, the quadratic
has 2 unequal Real solutions.
Cheers,
Stan H.